OpenStudy (anonymous):

Can any body tell me how to write 16 in binary? Thank you

OpenStudy (across):

\(16=2^4\longrightarrow10000\)

OpenStudy (unklerhaukus):

1=1 2=10 3=11 4=100 5=101 6=110 7=111 8=1000 9=1001 10=1010 11=1011 12=1100 13=1101 14=1110 15=1111 16=

Parth (parthkohli):

\[16 = 1\cdot 2^4 + 0\cdot 2^3 + 0 \cdot 2^2 + 0\cdot 2 + 0 \cdot 2^0 \implies 10000 \]